OpenStudy (anonymous):

find all zeros of the function x^2+10x+23

OpenStudy (anonymous):

\[x1=(-10+\sqrt{8})/2=-3.586; x2=(-10-\sqrt{8})/2=-6.414\]

mathslover (mathslover):

\[\huge{x^2+10x+23}\] Let this equation be p(x) so p(x)=0 \[\huge{x^2+10x+23=0}\] \[\huge{ax^2+bx+c=0}\] --- Just compare these equations that is 1) p(x) and 2) general form of a quadratic equation : The roots of a quadratic equation : \[\huge{x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}}\] put a = 1 , b = 10 , c = 23

mathslover (mathslover):

can you do this now @dorkkk

OpenStudy (anonymous):